OpenStudy (anonymous):

Tell whether the system has one solution, infinitely solution, or no solution. -x + y= -13 3x - y =19

OpenStudy (anonymous):

Notice that the y coefficiants cancel out very neatly =D So add the two equations like so: -x + y = -13 + 3x - y = 19 -------------- 2x = 6 now solve that for x. Plug in for x in either of the starting equations to solve for y. You get one answer.

OpenStudy (anonymous):